OpenStudy (anonymous):

a grocery store makes a breakfast cereal by mixing oat cereal costing $2 per kilogram with dried fruits costing $9 per kilogram. how many kilograms of dried fruit are needed to make 60 kg of cereal with a cost of $3.75 per kilogram?

OpenStudy (anonymous):

\[2x+9(60-x) = 37.5\times 60\]

OpenStudy (anonymous):

of course i meant \[=3.75 \times 60\]